OpenStudy (anonymous):

Using complete sentences, explain each step in simplifying the ratio x^2-x-20 over x^2+8x+16

OpenStudy (anonymous):

@itsmylife help again?c:

OpenStudy (anonymous):

what is it you dont understand?

OpenStudy (anonymous):

\[\frac{ x^2-x-20 }{ x^2+8x+16 }=\frac{ x^2-5x-4x-20 }{ x^2+4x+4x+16 }\]babe you can do it now ;)

OpenStudy (anonymous):

im still confused :c @itsmylife

OpenStudy (anonymous):

babe its simple \[\frac{ (x-5)(x+4) }{ (x+4) (x+4)}=\frac{ x-5 }{ x+4 }\]

OpenStudy (anonymous):

thers mistake in factorization in the neumrator it should be x^2-5x+4x-20 and np ;)

OpenStudy (anonymous):

help me with one more pleaseeeeee (: @itsmylife

OpenStudy (anonymous):

help me with one more pleaseeeeee (: @itsmylife

OpenStudy (anonymous):

whers it?

OpenStudy (anonymous):

one sec (: